# Partial sampling of a random spanning tree

### v3

@@ 1 Introduction
The theorem given below provides an answer in a slightly more general situation: One of the specific vertices denoted by $\Delta$ is chosen to be the root of $\Upsilon$. The weight of a tree is assumed to be proportional to $\kappa^{p}$, $\kappa$ being a positive parameter and $p$ the degree of $\Delta$ in the spanning tree. For $\kappa\neq 1$, the distribution we get on spanning tree is not entirely uniform.
@@ 2 Definitions, first properties and statement
$\det(G^{(n),\kappa}=\frac{1}{\kappa(n+\kappa)^{n-1}}$ (1)
@@
Coming back to the general case, a simple way to sample a spanning tree $\Upsilon$ under the probability $\mathbb{P}$ is to perform Wilson’s algorithm (cf. [12]) which is based on loop erasure. The probability of a loop-erased path from a given vertex $\xi_{0}$ to $\Delta$ is the product of its conductances normalized by the determinant of the restriction of the Green matrix to the vertices of $\xi$. Wilson’s algorithm starts by choosing any order on the vertices and constructing a loop erased path from the first vertex to $\Delta$, then at each step it constructs a loop erased path from the first unused vertex to the tree of used vertices (including $\Delta$) until all vertices have been used.
We are interested in the smallest subtree $\Upsilon_{L}$ of $\Upsilon$ connecting a set $L$ of $l$ vertices and $\Delta$. For that, we can run Wilson algorithm after choosing an order in which these vertices are the first $l$ vertices. Given a tree $Y$ imbedded in the graph and rooted in $\Delta$ containing $L$ and whose set of leaves is included in $L$ , the spanning tree $\Upsilon$ produced by the algorithm will contain it iff before the $l$-th step of the algorithm the tree of used vertices is exactly $Y$. Then we will have $\Upsilon_{L}=Y$.
It is easy to check from the proof of Wilson’s algorithm given in [8] that for any tree $Y$ rooted in $\Delta$, the probability $\mathbb{P}(Y=\Upsilon_{L})$ is given by the product of the conductances of the edges of $Y$ normalized by the determinant of the restriction of the Green matrix to the vertices of $Y$.
Consequently, in the case of the complete graph with $n$ vertices $\{1,...,n\}$ endowed with unit conductances and a constant killing factor $\kappa$ if $Y$ has $d$ vertices (root excepted) and $r$ edges incident to $\Delta$, by formula 2, we have:
@@
$\mathbb{P}^{(n),\kappa}(Y=\Upsilon_{L})=\frac{\kappa^{r-1}(d+\kappa)}{(n+\kappa)^{d}}$
@@
We say that a tree is $L$-reduced if all its vertices which are not elements of $L$ or nodes have been removed. A $L$-tree can be decomposed into its $L$-reduction which is still a $L$ tree and finite sequences of intermediate inner vertices with a single child.
It is clearly enough to consider the case $L=\{1,...l\}$. The order on $L$ is used to induce an order on subsets of $L$ by listing the leaves in the set in increasing order and then use alphabetical order. For example we have $1<(1,3,7)<2<(2,3)$. Nodes and leaves of a reduced $L$-tree inherit of that order we will call the $L$-order.
@@
We say that a $L$ tree is binary iff $L^{*}=L$, every internal vertex has exactly two children and the root only one. An easy induction shows that the set $\mathcal{B}_{l}$ of binary $L$-tree classes has cardinality $c_{l}=\prod_{i=1}^{l-1}(2i-1)=\frac{(2l-3)!}{2^{l-2}(l-2)!}$ and that these trees have l1l-1 internal vertices. Indeed, adding one leaf is done by choosing a vertex, add a node just below it and connect it to the new leaf. Consider now binary tree extensions: using the $L$-order on leaves and nodes, $u_{L}$ becomes a (2l1)(2l-1)-tuple of nonnegative integers $(u_{L}(i),1\leq i\leq 2l-1)$.
@@ 3 Proof
Then we start with the proof of c). Note that setting $\Sigma=\sum_{1}^{2l-1}[t_{i}\sqrt{n}]+l-1$, there are $\prod_{i=l}^{\Sigma+l}(n-i)$ possible choices for the inner vertices of $\Upsilon_{L}$, given that $u_{L}=[t\sqrt{n}]$. From lemma 2.1, we get that
@@
$\ln(\mathbb{P}^{(n),\kappa}(Q_{L}=\alpha,u_{L}(i)=[t_{i}\sqrt{n}]))=\ln(\Sigma+l+\kappa)-(\Sigma+l)\ln(n+\kappa)+(n-l)\ln(n-l)-n+l+\frac{1}{2}\ln(2\pi(n-l))-(n-\Sigma-l)\ln(n-\Sigma-l)+n-\Sigma-l-\frac{1}{2}\ln(2\pi(n-\Sigma-l))+O(1/n)$
which, after decomposing the first term into $1/2\ln(n)+\ln(\sum t_{i})+\ln((\Sigma+l+\kappa)/(\sum t_{i}\sqrt{n}))$ and each term of the form $\ln(n+a)$ into $\ln(n)+\ln(1+a/n)$, by Taylor’s formula applied to the $\ln$ functions gives:
$\ln(\mathbb{P}^{(n),\kappa}(Q_{L}=\alpha,u_{L}(i)=[t_{i}\sqrt{n}])=1/2\ln(n)+\ln(\sum t_{i})-l\ln(n)-\Sigma+n\ln(1-\frac{\Sigma-l}{n})+O(n^{-1/2})=\ln(\Sigma/\sqrt{n})-(2l-1)\ln(\sqrt{n})-\Sigma^{2}/2n+O(n^{-1/2}).$
@@ 4 Additional comments
6) As $l$ increases, we get from b) a consistent family of distributions on $\mathbb{R}_{+}$-extensions of binary tree classes with $l$ leaves. This family is the consistent family of « proper $k$-trees » defined by Aldous in section 4.3 of [3]. It is shown in [3] that this family can be represented by a « Brownian continuum random tree »0 (CRT) as it satisfies a tightness condition. Other constructions of this CRT are given in [2], [1], using Brownian excursions or branching processes. See also [9], [10], [6].

### v4

@@ 1 Introduction
The theorem given below provides an answer in a slightly more general situation: One of the specific vertices denoted by $\Delta$ is chosen to be the root of $\Upsilon$. The weight of a tree is assumed to be proportional to $\kappa^{p_{\Delta}}$, $\kappa$ being a positive parameter and $p_{\Delta}$ the degree of $\Delta$ in the tree. For $\kappa\neq 1$, the distribution we get on spanning tree is not entirely uniform.
@@ 2 Definitions, first properties and statement
$\det(G^{(n),\kappa})=\frac{1}{\kappa(n+\kappa)^{n-1}}$ (1)
@@
Coming back to the general case, a simple way to sample a spanning tree $\Upsilon$ under the probability $\mathbb{P}$ is to perform Wilson’s algorithm (cf. [12]) which is based on loop erasure. The probability of a loop-erased path $\xi$ from a given vertex $\xi_{0}$ to $\Delta$ is the product of its conductances multiplied by the determinant of the restriction of the Green matrix to the vertices of $\xi$. Wilson’s algorithm starts by choosing any order on the vertices and constructing a loop erased path from the first vertex to $\Delta$, then at each step it constructs a loop erased path from the first unused vertex to the tree of used vertices (including $\Delta$) until all vertices have been used.
We are interested in the smallest subtree $\Upsilon_{L}$ of $\Upsilon$ connecting a set $L$ of $l$ vertices and $\Delta$. For that, we can run Wilson algorithm after choosing an order in which these vertices are the first $l$ vertices. Given a tree $Y$ imbedded in the graph and rooted in $\Delta$ containing $L$ and whose set of leaves is included in $L$ , the spanning tree $\Upsilon$ produced by the algorithm will contain it iff after the $l$-th step of the algorithm the tree of used vertices is exactly $Y$. Then we will have $\Upsilon_{L}=Y$.
It is easy to check from the proof of Wilson’s algorithm given in [8] that for any tree $Y$ rooted in $\Delta$, the probability $\mathbb{P}(Y=\Upsilon_{L})$ is given by the product of the conductances of the edges of $Y$ multiplied by the determinant of the restriction of the Green matrix to the vertices of $Y\setminus\{\Delta\}$.
Consequently, in the case of the complete graph with $n$ vertices $\{1,...,n\}$ endowed with unit conductances and a constant killing factor $\kappa$ if $Y$ has $a$ vertices (root excepted) and $r$ edges incident to $\Delta$, by formula 2, we have:
@@
$\mathbb{P}^{(n),\kappa}(Y=\Upsilon_{L})=\frac{\kappa^{r-1}(a+\kappa)}{(n+\kappa)^{d}}$
@@
We say that a tree is $L$-reduced if all its vertices which are not elements of $L$ or nodes have been removed. A $L$-tree can be decomposed into its $L$-reduction which is still a $L$ -tree and finite sequences of intermediate inner vertices with a single child.
It is clearly enough to consider the case $L=\{1,...l\}$. Then the natural order on $L$ is used to induce an order on subsets of $L$ by listing the leaves in each set in increasing order and then use alphabetical order to compare their lists of increasing digits. For example we have $(1)<(1,2)<(1,3,7)<(2)<(2,3)$ etc.. Nodes and leaves of a reduced $L$-tree inherit of that order we will call the $L$-order.
@@
We say that a $L$-tree is binary iff $L^{*}=L$, every internal vertex has exactly two children and the root only one. An easy induction shows that the set $\mathcal{B}_{l}$ of binary $L$-tree classes has cardinality $c_{l}=\prod_{i=1}^{l-1}(2i-1)=\frac{(2l-3)!}{2^{l-2}(l-2)!}$ and that these trees have l1l-1 internal vertices. Indeed, adding one leaf is done by choosing a vertex, adding a node just below it and connecting it to the new leaf. Consider now binary tree extensions: using the $L$-order on leaves and nodes, $u_{L}$ becomes a (2l1)(2l-1)-tuple of nonnegative integers $(u_{L}(i),1\leq i\leq 2l-1)$.
@@ 3 Proof
Then we start with the proof of c). Note that setting $\Sigma=\sum_{1}^{2l-1}[t_{i}\sqrt{n}]$ there are $\prod_{i=l}^{\Sigma+l}(n-i)$ possible choices for the inner vertices of $\Upsilon_{L}$, given that $u_{L}=[t\sqrt{n}]$. From lemma 2.1, we get that
@@
$\ln(\mathbb{P}^{(n),\kappa}(Q_{L}=\alpha,u_{L}=[t\sqrt{n}]))=\ln(\Sigma+l+\kappa)-(\Sigma+l)\ln(n+\kappa)+(n-l)\ln(n-l)-n+l\\ +\frac{1}{2}\ln(2\pi(n-l))-(n-\Sigma-l)\ln(n-\Sigma-l)+n-\Sigma-l-\frac{1}{2}\ln(2\pi(n-\Sigma-l))+O(1/n)$ (3)
After decomposing the first term into $1/2\ln(n)+\ln(\Sigma/\sqrt{n})+\ln(1+(l+\kappa)/\Sigma)$ and each term of the form $\ln(n+a)$ into $\ln(n)+\ln(1+a/n)$, Taylor’s formula applied to the $\ln$ functions gives:
$\ln(\mathbb{P}^{(n),\kappa}(Q_{L}=\alpha,u_{L}(i)=[t_{i}\sqrt{n}])=1/2\ln(n)+\ln(\Sigma/\sqrt{n})-l\ln(n)-\Sigma\\ +n\ln(1-\frac{l}{n})-n\ln(1-\frac{\Sigma+l}{n})+O(n^{-1/2})=\ln(\Sigma/\sqrt{n})-(2l-1)\ln(\sqrt{n})-\Sigma^{2}/2n+O(n^{-1/2}).$ (4)
@@ 4 Additional comments
6) As $l$ increases, we get from b) a consistent family of distributions on $\mathbb{R}_{+}$-extensions of binary tree classes with $l$ leaves. This family is the consistent family of « proper $k$-trees » defined by Aldous in section 4.3 of [3]. It is shown in [3] that this family can be represented by a « Brownian continuum random tree » (CRT) as it satisfies a tightness condition. Other constructions of this CRT are given in [2], [1], using Brownian excursions or branching processes. See also [9], [10], [6].
7)It would be possible to let the parameter $\kappa$ depend on $n$, which might yield different asymptotic results.